OpenStudy (fluttershyk):

Hope nobody minds helping me out here LIKE REALLY HELP Ok so I've been in the hospital and haven't had time to get anything done so heres the question You've probably heard of Albert Einstein's formula regarding the speed of light E= mc2 while this is an important formula that you will use throughtout your academic career, it is not a formula you apply in your daily life. Describe some common or unique formulas that you use in your life.

OpenStudy (anonymous):

\[\frac{ -b \sqrt{b^2+4ac} }{ 2a }\]

OpenStudy (anonymous):

-4ac***

OpenStudy (fluttershyk):

@unthinkablejoe it didn't come through what ever you wrote

OpenStudy (fluttershyk):

@KamiBug help me please your the best at this stuff

OpenStudy (fluttershyk):

@truwhovian11 help me

OpenStudy (anonymous):

You get an allowance of $10 a week. You understand that the amount of money you have after x weeks will be equal to $10 times the number of weeks. *Expressed algebraically, this means: $ = weeks * $10 /week **Even more simply put : $ = 10 * w Understanding this equation, that $ = weeks * $10 /week, means we can see that after 4 weeks, our $ = 40. This is cool, because if you want something that is $ = 80 : 80 = 10 * w w = 80 /10 = 8 weeks Then you understand that it will take you 8 weeks to save enough allowance to get what you want.

OpenStudy (fluttershyk):

@UC66 thank you that's good I mean I didn't want a book but that's ok thank you

OpenStudy (fluttershyk):

@UC66 can you help me with some other multiple choice I mean like check the answers that I have and tell me if im right wrong or close